# Predicting random close packing of binary hard-disk mixtures via third-virial-based parameters

### v1

@@ Abstract
We propose a simple and accurate approach to estimate the random close packing (RCP) fraction of binary hard-disk mixtures. By introducing a parameter based on the reduced third virial coefficient of the mixture, we show that the RCP fraction depends nearly linearly on this parameter, leading to a universal collapse of simulation data across a wide range of size ratios and compositions. Comparisons with previous models by Brouwers and Zaccone demonstrate that our approach provides the most consistent and accurate predictions. The method can be naturally extended to polydisperse mixtures with continuous size distributions, offering a robust framework for understanding the universality of RCP in hard-disk systems.
@@ I Introduction
Ever since the introduction of the concept by Bernal [1, 2, 3, 4, 5], random close packing (RCP) has remained a subject of sustained interest across mathematics, physics, and engineering [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. This enduring attention stems from the ubiquitous appearance of densely packed disordered structures in a wide variety of systems, including jamming transitions, colloids, granular media, glasses, amorphous solids, ceramics, and even living matter, as illustrated by the extensive body of work cited above.
A commonly adopted definition of RCP—although sometimes challenged as ill-defined—identifies it as the highest achievable density at which a large collection of particles can be packed without long-range order. The concept was originally formulated for monodisperse hard spheres and later extended to monodisperse hard disks, systems that have played a paradigmatic role in the study of packing problems. An additional complication arises from the fact that different packing protocols, whether numerical or experimental, may lead to different maximal densities [27, 37].
@@
In two recent papers [63, 64], Brouwers proposed, using a geometric approach, the following approximation for the RCP fraction ($\phi_{\text{mixt}}$) of a binary HD mixture with small or moderate dispersity:
@@
Here, $\phi_{\text{mono}}$ denotes the RCP fraction of the monodisperse HD system, $x_{\text{L}}$ and $x_{\text{S}}=1-x_{\text{L}}$ are the mole fractions of large and small disks, respectively, $q\equiv\sigma_{\text{L}}/\sigma_{\text{S}}\geq 1$ is the size ratio, and $m_{n}=M_{n}/M_{1}^{n}$ are reduced moments, with $M_{n}=x_{\text{S}}\sigma_{\text{S}}^{n}+x_{\text{L}}\sigma_{\text{L}}^{n}$ the moments of the size distribution.
By fitting Eq. (1a) to Monte Carlo simulation data [73] for $q=1.4$, Brouwers obtained $\phi_{\text{mono}}=0.8425$, a value slightly below the commonly accepted estimate $\phi_{\text{mono}}=0.844$ [37].
Figure 1: (a) $1/\phi_{\text{mixt}}$ versus the parameter $\mu_{B}$ [Eq. (1a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda_{B}$ [Eq. (1b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73], while the dashed and dashed-dotted lines correspond to Eqs. (1a) and (1b) with $\phi_{\text{mono}}=0.844$ and $\phi_{\text{mono}}=0.8425$, respectively.
Two noteworthy consequences follow from Eqs. (1). First, the entire dependence of $\phi_{\text{mixt}}$ on $x_{\text{L}}$ and $q$ is encoded in the single parameter $\mu_{B}$ (or, equivalently, in $\lambda_{B}$). Consequently, two distinct binary mixtures sharing the same value of $\mu_{B}$ are predicted to have the same RCP fraction. Second, the inverse RCP fraction $1/\phi_{\text{mixt}}$ depends linearly on $\mu_{B}$, while the ratio $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda_{B}$. As shown in the Appendix, these two linear dependencies uniquely fix the functional form of Eqs. (1). This implies that plots of $1/\phi_{\text{mixt}}$ versus $\mu_{B}$ and $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda_{B}$ for different mixtures should exhibit data collapse.
@@
The main purpose of this short paper is to demonstrate that a significantly better collapse is achieved by slightly modifying Eqs. (1) and introducing an alternative “universality” parameter based on the third virial coefficient of the mixture. We further demonstrate that this approach can be naturally extended to polydisperse mixtures, providing a universal and practically useful framework for predicting RCP fractions in both discrete and continuous size distributions.
Figure 2: (a) $\phi_{\text{mixt}}$ versus the parameter $\mu$ [Eq. (2a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda$ [Eq. (2b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73], and the dashed lines correspond to our proposal [Eqs. (2)] with $\phi_{\text{mono}}=0.844$.
Figure 3: $\phi_{\text{mixt}}$ as a function of the area fraction of large disks, $c_{\text{L}}$, for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73]. Solid lines show our predictions [Eq. (2a)] with $\phi_{\text{mono}}=0.844$, while the dashed-dotted and dashed lines correspond to Eq. (1a) with $\phi_{\text{mono}}=0.8425$ and Eq. (II) with $\phi_{\text{mono}}=0.844$, respectively.
@@ II Our proposal
In our approach, we replace Eqs. (1) with the following expressions:
@@
where $b_{3}=4\left(\frac{4}{3}-\frac{\sqrt{3}}{\pi}\right)\simeq 3.12802$, and
$\bar{B}_{3}=$ $\frac{1}{m_{2}^{2}}\left[x_{\text{S}}^{3}b_{3}+\frac{x_{\text{S}}^{2}x_{\text{L}}}{\pi}F(q)+\frac{x_{\text{S}}x_{\text{L}}^{2}}{\pi}q^{4}F(q^{-1})+x_{\text{L}}^{3}q^{4}b_{3}\right]$ (3)
is the reduced third virial coefficient of the binary mixture [74], with
@@
The parameters $\mu$ and $\lambda$ defined in Eqs. (2) are motivated by the “surplus” approximation for the equation of state of HD or HS fluid mixtures [75, 48, 74, 76].
According to Eqs. (2), $\phi_{\text{mixt}}$ depends linearly on $\mu$, while $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda$. As discussed in the Appendix, these two assumptions again uniquely determine the structure of Eqs. (2). Figure 2 shows that these simple linear relationships are well satisfied with the standard value $\phi_{\text{mono}}=0.844$, particularly for $q=3$, and that the collapse of data from different mixtures is significantly improved compared with Brouwers’ model.
A complementary proposal for HD mixtures has recently been put forward by Zaccone [68, 69]. In this approach, the RCP fraction is determined by equating the compressibility factors of the mixture and the monodisperse system, $Z_{\text{mixt}}(\phi_{\text{mixt}})=Z_{\text{mono}}(\phi_{\text{mono}})$. Using the Scaled Particle Theory expressions [77, 78, 79], this leads to the explicit formula [69]:
@@
for $q=1.4$, $1.7$, 22, and $3$. As observed, Brouwers’ proposal [Eq. (1a)] with the fitted value $\phi_{\text{mono}}=0.8425$ is accurate for $q=1.4$ and reasonably good for $q=1.7$, by construction, but fails for $q=2$ and $q=3$. Zaccone’s approach [Eq. (II)] with $\phi_{\text{mono}}=0.844$ performs fairly well for $q=1.4$, but underestimates $\phi_{\text{mixt}}$ for larger size ratios. Using the monodisperse value $\phi_{\text{mono}}\simeq 0.886$ [80] in Eq. (II), as done in Refs. [68, 69], leads to even poorer estimates. In contrast, our third-virial-based proposal provides the best agreement across all size ratios, demonstrating its effectiveness in capturing the universality of RCP in binary HD mixtures.
@@ III Extension to polydisperse mixtures
Figure 4: Plot of the parameter $\mu$ as a function of the dispersity parameter $s$ for a polydisperse mixture with a log-normal size distribution.
We conjecture that Eq. (2a) remains a good candidate for estimating the RCP fraction of a polydisperse HD mixture characterized by a continuous size distribution $f(\sigma)$. In this case, the moments of the distribution are defined as $M_{n}=\int_{0}^{\infty}d\sigma,f(\sigma)\sigma^{n}$, and the reduced third virial coefficient is given by [74]
@@
$\times\sigma_{12}^{2}\,\mathcal{S}_{\sigma_{13},\sigma_{23}}(\sigma_{12}),$ (7)
where
@@
$-\frac{1}{2}\sqrt{2r^{2}(a^{2}+b^{2})-(b^{2}-a^{2})^{2}-r^{4}}.$ (8)
is the intersection area of two circles of radii $a$ and $b$ whose centers are separated by a distance $r$.
@@
for which the moments are $M_{n}=\sigma_{0}^{n}e^{n^{2}s^{2}/2}$ and the reduced moments $m_{n}=e^{n(n-1)s^{2}/2}$. The value of $\bar{B}_{3}$ must then be evaluated numerically via the triple integral in Eq. (7). Figure 4 shows the parameter $\mu$ as a function of the dispersity parameter $s$. According to Eq. (2a), the RCP fraction $\phi_{\text{mixt}}$ is expected to vary approximately linearly with $\mu(s)$.
In general, the values of $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ may depend on the packing protocol employed [27, 37]. Nevertheless, Eqs. (2) are expected to remain applicable provided that $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ are determined using equivalent protocols.
@@ IV Conclusions
In this work, we have proposed a simple yet effective modification of Brouwers’ geometric approach for estimating the RCP fraction of binary HD mixtures. By introducing a parameter based on the mixture’s reduced third virial coefficient, we obtained a nearly linear dependence of $\phi_{\text{mixt}}$ on $\mu$ and of $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ on $\lambda$, leading to a significantly improved collapse of simulation data across a wide range of size ratios. Comparisons with Brouwers’ original model and Zaccone’s proposal demonstrate that our approach provides the most accurate description of the RCP fraction for both small and moderate dispersity.
Furthermore, the method can be naturally extended to polydisperse mixtures with continuous size distributions, such as log-normal distributions, by computing the corresponding moments and third virial coefficient. The resulting parameter $\mu$ offers a universal measure of dispersity, allowing $\phi_{\text{mixt}}$ to be predicted as a nearly linear function of $\mu$.
Although this extension to polydisperse mixtures remains a conjecture, the excellent agreement with simulation results in the binary case suggests that it is worth testing. Its simplicity and predictive potential may stimulate new simulations to either confirm or refute the conjecture, providing a clear avenue for future research.
Overall, our results suggest that the third-virial-based parameter provides a robust and practically useful framework for understanding the universality of RCP in HD mixtures, with potential applications to more complex and polydisperse systems.

### v3

@@ Abstract
We propose a simple and accurate approach to estimate the random close packing (RCP) fraction of binary hard-disk mixtures. By introducing a parameter based on the mixture’s reduced third virial coefficient—which effectively captures three-body correlations and excluded-area constraints—we show that the RCP fraction depends nearly linearly on this parameter, leading to a near-universal collapse of simulation data over a wide range of size ratios and compositions. Comparisons with previous models by Brouwers and Zaccone indicate that the present approach provides more accurate and consistent predictions. The method can be naturally extended to polydisperse mixtures with continuous size distributions and is structurally consistent with the surplus equation-of-state formulation, offering a compact framework for understanding the near universality of RCP in hard-disk systems.
@@ I Introduction
Ever since the introduction of the concept by Bernal,[1, 2, 3, 4, 5] random close packing (RCP) has remained a subject of sustained interest across mathematics, physics, and engineering.[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] This enduring attention stems from the ubiquitous appearance of densely packed disordered structures in a wide variety of systems, including jamming transitions,[4, 28, 30, 31, 38, 39, 40, 43, 50, 51, 52, 53, 56, 66] colloids,[10, 12, 13] granular media,[14, 46, 49] glasses,[42] amorphous solids,[18] ceramics,[57] and even living matter,[76] as illustrated by the extensive body of work cited above.
A commonly adopted definition of RCP, although sometimes challenged as ill-defined,[27, 34] identifies it as the highest achievable density at which a large collection of particles can be packed without long-range order.[17, 34, 54] The concept was originally formulated for monodisperse hard-sphere (HS) systems and later extended to monodisperse hard-disk (HD) systems, both of which have played a paradigmatic role in the study of packing problems. An additional complication arises from the fact that different packing protocols, whether numerical or experimental, may lead to different maximal densities.[27, 37]
@@
In two recent papers,[67, 68] Brouwers proposed, using a geometric approach, the following approximation for the RCP fraction ($\phi_{\text{mixt}}$) of a binary HD mixture with small or moderate dispersity:
@@
Here, $\phi_{\text{mono}}$ denotes the RCP fraction of the monodisperse HD system, $x_{\text{L}}$ and $x_{\text{S}}=1-x_{\text{L}}$ are the mole fractions of large and small disks, respectively, $q\equiv\sigma_{\text{L}}/\sigma_{\text{S}}\geq 1$ is the size ratio, and $m_{n}=M_{n}/M_{1}^{n}$ are reduced moments, with $M_{n}=x_{\text{S}}\sigma_{\text{S}}^{n}+x_{\text{L}}\sigma_{\text{L}}^{n}$ being the moments of the size distribution.
By fitting Eq. (1a) to Monte Carlo simulation data for q=1.4q=1.4,[78] Brouwers obtained $\phi_{\text{mono}}=0.8425$, a value slightly below the commonly accepted estimate $\phi_{\text{mono}}=0.844$.[37]
Figure 1: (a) $1/\phi_{\text{mixt}}$ vs the parameter $\mu_{B}$ [Eq. (1a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ vs $\lambda_{B}$ [Eq. (1b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data,[78] while the dashed and dashed-dotted lines correspond to Eqs. (1a) and (1b) with $\phi_{\text{mono}}=0.844$ and $\phi_{\text{mono}}=0.8425$, respectively.
Two noteworthy consequences follow from Eq. (1). First, the entire dependence of $\phi_{\text{mixt}}$ on $x_{\text{L}}$ and $q$ is encoded in the single parameter $\mu_{B}$ (or, equivalently, in $\lambda_{B}$). Consequently, two distinct binary mixtures sharing the same value of $\mu_{B}$ are predicted to have the same RCP fraction. Second, the inverse RCP fraction $1/\phi_{\text{mixt}}$ depends linearly on $\mu_{B}$, while the ratio $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda_{B}$. As shown in Appendix A, these two linear dependencies uniquely fix the functional form of Eq. (1). This implies that plots of $1/\phi_{\text{mixt}}$ vs $\mu_{B}$ and $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ vs $\lambda_{B}$ for different mixtures should exhibit data collapse.
@@
The main purpose of this short paper is to demonstrate that a significantly better collapse is achieved by slightly modifying Eq. (1) and introducing an alternative “universality” parameter based on the third virial coefficient of the mixture. We further demonstrate that this approach can be naturally extended to polydisperse mixtures, providing a universal and practically useful framework for predicting RCP fractions in both discrete and continuous size distributions.
Figure 2: (a) $\phi_{\text{mixt}}$ vs the parameter $\mu$ [Eq. (2a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ vs $\lambda$ [Eq. (2b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data,[78] and the dashed lines correspond to our proposal [Eq. (2)] with $\phi_{\text{mono}}=0.844$.
Figure 3: $\phi_{\text{mixt}}$ as a function of the area fraction of large disks, $c_{\text{L}}$, for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data.[78] Solid lines show our predictions [Eq. (2a)] with $\phi_{\text{mono}}=0.844$, while the dashed-dotted and dashed lines correspond to Eq. (1a) with $\phi_{\text{mono}}=0.8425$ and Eq. (II) with $\phi_{\text{mono}}=0.844$, respectively.
@@ II Our proposal
In our approach, we replace Eq. (1) with the following expressions:
@@
where $b_{3}=4\left(\frac{4}{3}-\frac{\sqrt{3}}{\pi}\right)\simeq 3.128\thinspace 02$, and
$\bar{B}_{3}=$ $\frac{\sigma_{\text{S}}^{4}}{M_{2}^{2}}\left[x_{\text{S}}^{3}b_{3}+\frac{x_{\text{S}}^{2}x_{\text{L}}}{\pi}F(q)+\frac{x_{\text{S}}x_{\text{L}}^{2}}{\pi}q^{4}F(q^{-1})+x_{\text{L}}^{3}q^{4}b_{3}\right]$ (3)
is the reduced third virial coefficient of the binary mixture,[79] with
@@
The parameters $\mu$ and $\lambda$ defined in Eq. (2) are motivated by the “surplus” approximation for the equation of state of $d$-dimensional fluid mixtures,[80, 48, 79, 81, 82] as detailed in Appendix B.
The key physical idea underlying Eq. (2) is that three-body correlations, encoded in the reduced third virial coefficient $\bar{B}_{3}$, capture the dominant effects of size disparity on dense disordered packing. In this picture, local triplet constraints and excluded-area effects largely determine how efficiently disks can arrange at jamming. Related viewpoints have recently been employed[56] to estimate the RCP density of three-dimensional monodisperse spherical and nonspherical particles by identifying it with the singular point obtained from the analytical continuation of the liquid-branch equation of state.
According to Eq. (2), $\phi_{\text{mixt}}$ depends linearly on $\mu$, while $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda$. As discussed in Appendix A, these two assumptions again uniquely determine the structure of Eq. (2). Figure 2 shows that these simple linear relationships are well satisfied with the standard value $\phi_{\text{mono}}=0.844$, particularly for $q=3$, and that the collapse of data from different mixtures is significantly improved compared with Brouwers’ model.
A complementary proposal for HD mixtures has recently been put forward by Zaccone.[72, 73] In this approach, the RCP fraction is determined by equating the compressibility factors of the mixture and the monodisperse system, $Z_{\text{mixt}}(\phi_{\text{mixt}})=Z_{\text{mono}}(\phi_{\text{mono}})$. Using the Scaled Particle Theory expressions,[83, 84, 85] this leads to the explicit formula:[73]
@@
for $q=1.4$, $1.7$, 22, and $3$. As observed, Brouwers’ proposal [Eq. (1a)] with the fitted value $\phi_{\text{mono}}=0.8425$ is accurate for $q=1.4$ and reasonably good for $q=1.7$, by construction, but it fails for $q=2$ and $q=3$. Zaccone’s approach [Eq. (II)] with $\phi_{\text{mono}}=0.844$ performs fairly well for $q=1.4$, but it underestimates $\phi_{\text{mixt}}$ for larger size ratios. Using the monodisperse value $\phi_{\text{mono}}\simeq 0.886$[86] in Eq. (II), as done in Refs. 72, 73, leads to even poorer estimates. In contrast, our third-virial-based proposal provides the best agreement across all size ratios, demonstrating its effectiveness in capturing the universality of RCP in binary HD mixtures.
@@ III Extension to polydisperse mixtures
Figure 4: Plot of the parameter $\mu$ as a function of the dispersity parameter $s$ for a polydisperse mixture with a log-normal size distribution. The inset shows a logarithmic plot of $1-\mu$ as a function of $s^{2}$. The dashed-dotted line has a slope of 0.4-0.4.
We conjecture that Eq. (2a) remains a good candidate for estimating the RCP fraction of a polydisperse HD mixture characterized by a continuous size distribution $f(\sigma)$. In this case, the moments of the distribution are defined as $M_{n}=\int_{0}^{\infty}d\sigma\,f(\sigma)\sigma^{n}$, and the reduced third virial coefficient is given by[79]
@@
$\times{A}(\sigma_{1},\sigma_{2},\sigma_{3}),$ (7)
where ${A}(\sigma_{1},\sigma_{2},\sigma_{3})\equiv\sigma_{12}^{2}\mathcal{S}_{\sigma_{13},\sigma_{23}}(\sigma_{12})$ and
@@
$-\frac{1}{2}\sqrt{2r^{2}(a^{2}+b^{2})-(b^{2}-a^{2})^{2}-r^{4}}$ (8)
is the intersection area of two circles of radii $a$ and $b$ whose centers are separated by a distance $r$. Note that ${A}(\sigma_{1},\sigma_{2},\sigma_{3})$ is a homogeneous function of degree four.
@@
for which the moments are $M_{n}=\sigma_{0}^{n}e^{n^{2}s^{2}/2}$ and the reduced moments are $m_{n}=e^{n(n-1)s^{2}/2}$. By introducing the new variables
$u=\frac{1}{3s}\ln\frac{\sigma_{1}\sigma_{2}\sigma_{3}}{\sigma_{0}^{3}},\quad y_{1}=\frac{1}{s}\ln\frac{\sigma_{1}}{\sigma_{0}}-u,\quad y_{2}=\frac{1}{s}\ln\frac{\sigma_{2}}{\sigma_{0}}-u,$ (10)
one has
$\bar{B}_{3}=$ $\frac{4\sqrt{3}e^{-4s^{2}/3}}{\pi^{2}}\int_{-\infty}^{\infty}dy_{1}\int_{-\infty}^{\infty}dy_{2}\,e^{-(y_{1}^{2}+y_{2}^{2}+y_{1}y_{2})}$
$\times{A}(e^{y_{1}s},e^{y_{2}s},e^{-(y_{1}+y_{2})s}),$ (11)
where we have taken into account the homogeneity property of the function $A(\sigma_{1},\sigma_{2},\sigma_{3})$ and performed the Gaussian integral over $u$. The double integral in Eq. (11) must be evaluated numerically.
Figure 4 shows the parameter $\mu$ as a function of the dispersity parameter $s$. According to Eq. (2a), the RCP fraction $\phi_{\text{mixt}}$ is expected to vary approximately linearly with $\mu(s)$. It ranges from $\mu=0$ as $s\to 0$ to $\mu=1$ as $s\to\infty$.
The asymptotic behavior $\mu\to 1^{-}$ for large $s$ ensures that Eq. (2a) remains consistent with the physical bound $\phi_{\text{mixt}}\leq 1$. The inset of Fig. 4 shows that $1-\mu$ decays approximately exponentially with $s^{2}$, indicating that the saturation is smooth rather than divergent. Thus, within the present third-virial-based framework, the mapping does not lead to unphysical packing fractions, even at large dispersity.
It should be stressed, however, that extremely large values of $s$ correspond to size distributions with exponentially broad tails, for which the physical interpretation of RCP itself becomes subtle. In such regimes, higher-order virial coefficients or a different functional form beyond the present linear mapping may be required to capture the packing constraints accurately. The present results suggest that the third virial coefficient captures the dominant geometric effects over a broad, but not arbitrarily extreme, range of dispersities.
The above considerations are independent of the specific choice of size distribution. In general, the values of $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ may depend on the packing protocol employed.[27, 37] Nevertheless, Eq. (2) is expected to remain applicable to arbitrary polydisperse mixtures, provided that $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ are determined using equivalent protocols.
@@ IV Conclusions
In this work, we have proposed a simple yet effective modification of Brouwers’ geometric approach for estimating the RCP fraction of binary HD mixtures. By introducing a parameter based on the mixture’s reduced third virial coefficient—which effectively encodes three-body correlations and excluded-area constraints—we obtained a linear dependence of $\phi_{\text{mixt}}$ on $\mu$ and of $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ on $\lambda$, leading to a significantly improved collapse of simulation data across a wide range of size ratios. Comparisons with Brouwers’ original model and Zaccone’s proposal show that our approach provides a more accurate description of the RCP fraction for both small and moderate dispersity.
Furthermore, the method can be naturally extended to polydisperse mixtures with continuous size distributions, such as log-normal distributions, by computing the corresponding moments and third virial coefficient. The resulting parameter $\mu$ offers a compact and universal measure of dispersity, allowing $\phi_{\text{mixt}}$ to be predicted as a nearly linear function of $\mu$.
Although this extension to continuous polydisperse mixtures remains a conjecture, the excellent agreement with simulation results in the binary case suggests that it is worth testing. Its simplicity and predictive character may stimulate new simulations aimed at confirming or refuting the conjecture, thereby providing a clear direction for future work.
An interesting open question concerns the possible extension of the present framework to mixtures of nonspherical convex particles, such as ellipses or spherocylinders. In such systems, the third virial coefficient incorporates not only size effects but also orientational excluded-area (or excluded-volume) contributions.[87, 88, 89, 90, 91, 92, 93] It is, therefore, tempting to speculate that a suitably defined reduced third virial coefficient, combined with the corresponding monocomponent RCP fraction for the given particle shape, could provide a natural generalization of the present linear mapping in terms of $\mu$. Whether this remains a quantitatively robust approximation in the presence of orientational degrees of freedom is, however, an issue that deserves dedicated investigation.
More generally, the present formulation is structurally consistent with the surplus equation-of-state framework for mixtures, reinforcing the idea that low-order virial information captures the dominant geometric constraints governing dense disordered packings. Overall, our results indicate that the third-virial-based parameter $\mu$ provides a robust and practically useful framework for understanding the near-universality of RCP in HD mixtures, with potential extensions to more complex and polydisperse systems.